International Year of Astronomy IYA 2009 From http://www.astronomy2009.org/ 2009 has been declared International Year of Astronomy by the General Assembly of the UN (United

Nations Organization).

The Opening ceremony of the International Year of Astronomy will take place in Paris, on January 15-16, 2009.

Catherine Cesarsky, is the president, since 2006, of the International Astronomical Union (IAU), which launched the International Year of Astronomy (2009). Calling it the golden age of astronomy, she said,

”Astronomy is the science that allows humanity to try to answer questions it has always asked. Where do we come from? Where are we going? Are we alone in the universe?”

”Our goal [in IYA] is to share with the rest of the world the wonder we feel as astronomers, faced with the mysteries of the universe. We want all countries in the world—many of which have already created programmes—to get involved. And we want the public, the widest possible public, to be able to take part. By the end of the year, we would like everyone on Earth to have spent at least a short while with their attention turned to the sky. Or at least to have read something on the most recent discoveries, or reflected on our position in the universe.

More information can be found on their web-site. Information on the Indian programmes are listed at the website http://www.iucaa.ernet.in/ iya09ind/ hosted by IUCAA, Pune.

As part of IYA, Jantar Mantar will carry some article(s) on astronomy in each issue this year. Look out for them in this and the following issues. Happy reading!